OpenStudy (anonymous):

Solve the equation: x^2 - 2x - 35+0

OpenStudy (raden):

x^2 - 2x - 35+0 (x-7)(x+5)=0 solve for x by take zeros of the factors

OpenStudy (raden):

i think ur equation : x^2 - 2x - 35=0, right ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

ok, now take each factors be zero it means : x-7=0 ----> x=...? x+5=0 ----> x=...?

OpenStudy (raden):

can u do it

OpenStudy (raden):

for x-7=0 add by 7 to both sides, what u get ?

OpenStudy (raden):

x-7+7=0+7 x=7 ok, do same idea for x+5=0

OpenStudy (anonymous):

using the t bar right

OpenStudy (raden):

what means " t bar " ?

OpenStudy (anonymous):